a steel plug has a diameter of 5cm at 30°c .atehat temperature will it fit exactly into a hole of constant diameter 4.997cm ( coefficient of liner expansion of stee is 11×10-60C-1

1
Expert's answer
2022-03-11T08:28:48-0500

The contraction of the diameter "\Delta d" of the plug with temperature is described as follows:


"\Delta d = \alpha d\Delta T = \alpha d(T_2-T_1)"

where "\Delta d = 4.997cm - 5cm = -0.003cm = -3\times 10^{-5}m", "\alpha = 11\times 10^{-6}\degree C^{-1}" is the coefficient of linear expansion of steel, "d = 5cm = 5\times 10^{-2}m" is the current diameter, "T_1 = 30\degree C" is the initial temperature, and "T_2" is the final required temperature. Solving for "T_2", obtain:


"T_2 = T_1 + \dfrac{\Delta d}{\alpha d}\\
T_2 = 30\degree C + \dfrac{-3\times 10^{-5}m}{11\times 10^{-6}\degree C^{-1}\cdot 5\times 10^{-2}} \approx -25\degree C"

Answer. "-25\degree C".